# Detecting low left ventricular ejection fraction from ECG using an interpretable and scalable predictor-driven framework

### 30 Mar 2026 v1

@@ 1 Introduction
Heart failure (HF) is a major global health burden, affecting over 64 million individuals worldwide, with prevalence continuing to rise (Savarese et al., 2022). In many patients, left ventricular systolic dysfunction (LVSD) is initially asymptomatic, yet it frequently progresses to symptomatic HF and is associated with poor prognosis if left undetected. Low left ventricular ejection fraction (LEF) is a critical indicator that reflects the progression of LVSD toward clinically manifest HF (Heidenreich et al., 2022). Early detection of LVSD is critical, as timely initiation of evidence-based therapies can mitigate symptoms, slow disease progression, and improve survival. Although cardiovascular imaging, particularly echocardiography (ECHO), remains the clinical gold standard for assessing ventricular function, it is time- and resource-intensive, requires specialized operators and equipment, and is not readily scalable for population-level screening (Ciampi and Villari, 2007). These limitations underscore the need for cost-effective, widely accessible approaches capable of reliably identifying patients with LEF (Attia et al., 2019; Yao et al., 2021; Tran et al., 2025; Lakshmanan and Mbanze, 2023).
@@
Electrocardiography (ECG), which records cardiac electrical activity, is a promising tool for large-scale screening (Hughes et al., 2024). Since its introduction by Willem Einthoven, which earned him the Nobel Prize in 1924, ECG science has steadily advanced: acquisition has become easier, signal quality has improved, and pathognomonic features for a wide range of rhythm and conduction disorders have been defined (Khera, 2024). More recently, artificial intelligence (AI) has transformed ECG analysis, enabling models to achieve cardiologist-level performance in traditional interpretation tasks (Hannun et al., 2019; Ribeiro et al., 2020; Jiang et al., 2024; Li et al., 2025) and to detect subtle patterns associated with conditions such as aortic stenosis (Kwon et al., 2020), tricuspid regurgitation (Diao et al., 2025), and LEF (Attia et al., 2019), which are often imperceptible to humans. Despite these advances, clinical adoption varies: AI-ECG models are widely accepted for well-established conditions such as atrial fibrillation and sinus bradycardia, which can be visually confirmed according to guidelines (Joglar et al., 2024; Kusumoto et al., 2019), but remain limited for emerging tasks such as LEF detection due to concerns about interpretability and the difficulty of linking predictions to recognizable ECG features (Van De Leur et al., 2022; Hughes et al., 2024). To enhance interpretability and facilitate broader clinical deployment, a recent study proposed a random forest model based on 555 discrete ECG measurements based on commercial algorithms, which demonstrated superior performance compared with N-terminal prohormone brain natriuretic peptide (NT-proBNP) for LEF detection (Hughes et al., 2024). Although this approach showed promising results, its performance is slightly below deep learning. Moreover, as highlighted in the original study, many features cannot be obtained when applied across different ECG machines, and measurement values differ markedly between devices (Strodthoff et al., 2023), which presents a significant barrier to broader clinical adoption.
@@
In this study, we aim to evaluate the efficacy of interpretable methods for detecting LEF using standard 12-lead ECGs. We propose ECG-based Predictor-Driven LEF (ECGPD-LEF), a framework that integrates traditional AI-ECG interpretation models with either a single-predictor approach or a multi-predictor approach to enhance recognition of subtle ECG features and improve LEF detection. The single-predictor approach performs zero-shot-like inference, requiring no additional training on the target dataset. The multi-predictor approach was trained on the publicly available ECG-ECHO paired dataset EchoNext, a comprehensive dataset encompassing diverse populations across races and clinical contexts. We also constructed an external validation dataset, MIMIC-LEF, based on MIMIC-IV (Johnson et al., 2024), MIMIC-IV-ECG (Gow et al., 2023), and MIMIC-IV-Note(Johnson et al., 2023b), which similarly covers a heterogeneous population. The model was evaluated on multi-center datasets, including internal validation (the EchoNext test set) and external validation (MIMIC-LEF). The model performance was assessed across multiple dimensions, including the presence or absence of other structural heart disease (SHD) and valvular heart disease (VHD), as well as across different care settings and racial or ethnic subgroups. Global and local model explanations were performed to identify important predictors contributing to LEF detection, and to characterize these subtle ECG changes associated with LEF.
@@ 2.1 Data sources
For the detection of LEF, we used the ECG-ECHO paired EchoNext dataset collected at Columbia University Irving Medical Center (Elias and Finer, 2025). External validation was performed using an ECG-Note paired dataset constructed from MIMIC-IV (Johnson et al., 2024), MIMIC-IV-ECG (Gow et al., 2023), and MIMIC-IV-Note (Johnson et al., 2023b), comprising records from Beth Israel Deaconess Medical Center.
@@ 2.3 ECG-ECHO and ECG-Note datasets
The ECG-Note dataset is constructed based on MIMIC-IV (Johnson et al., 2024), MIMIC-IV-ECG (Gow et al., 2023), and MIMIC-IV-Note (Johnson et al., 2023b), as illustrated in Figure A.1 of the Appendix. We initially identified adult patients who possessed at least one standard 10-second 12-lead ECG with sampling frequency of 500Hz and a corresponding clinical note recorded within one year following the ECG. Exclusion criteria were applied to remove: 1) pairs where clinical notes lacked keywords related to ejection fraction(EF) and ECHO/TTE; 2) ECGs containing missing data (NaN). Following these exclusions, the samples were cross-referenced with MIMIC-ECG-LVEF (Li et al., 2025) to identify consistent ECG-EF class pairs. For each patient, only the ECG-Note pair with the shortest time interval was retained, resulting in a final testing set of 16,017 ECGs.
@@ 2.7 Interpretability
For the multi-predictor approach, we employed SHAP (SHapley Additive exPlanations) values (Shapley and others, 1953; Lundberg and Lee, 2017) to quantify feature contributions at global and local levels. SHAP is a game-theoretic, additive feature attribution method that provides consistent and locally accurate explanations (Lundberg and Lee, 2017). For computational efficiency in tree-based models, we used the Tree SHAP algorithm (Lundberg et al., 2018). Global explanation plots included cumulative contributions, beeswarm summaries, and SHAP versus predictor value plots to identify key predictors and characterize their behavior in the model. Local plots highlighted the ten predictors SHAP values and displayed predicted probabilities relative to both the ECG diagnosis thresholds and the LEF decision thresholds, facilitating interpretation of individual model predictions.
@@ 3.1 Population characteristics
  • The training, validation, and internal test splits correspond to the official splits of the ECG-ECHO dataset, EchoNext (Elias and Finer, 2025). The external test cohort, ECG-Note, was derived from the MIMIC-IV database and its associated ECG and clinical note modules (Johnson et al., 2024; Gow et al., 2023; Johnson et al., 2023b). Values are shown as counts and percentages.
@@ 4.2 The need for publicly benchmarked AI-ECG evaluation
Artificial intelligence applied to ECG has demonstrated strong performance in emerging tasks, including detection of aortic stenosis, tricuspid regurgitation and left ventricular dysfunction. However, most prior models have been developed and evaluated on institution-specific private datasets with heterogeneous population characteristics and outcome definitions, limiting reproducibility and preventing rigorous cross-model comparison. Early work by Attia et al. (2019) trained a convolutional neural network on 44,959 patients to identify ventricular dysfunction defined as ejection fraction $\leq$ 35%, with prospective validation performed at the same institution (Attia and others, 2019). Subsequent refinements extended detection to ejection fraction $\leq$ 40% and reported multisite validation using digital ECG input alone (Carter et al., 2026). Although these studies achieved strong discrimination, the absence of publicly accessible benchmarking datasets constrains transparent evaluation. Recently, Poterucha et al. (2025) addressed this gap by releasing a de-identified ECG dataset comprising 36,286 unique patients with predefined training, validation, and test splits, and by benchmarking a baseline model (the Columbia mini model) that achieved discrimination comparable to models trained on substantially larger proprietary cohorts. By providing standardized data partitions, model weights, and open-source code, this work established a transparent and reproducible evaluation framework for AI-ECG research, upon which our study enables rigorous and fair comparison. Furthermore, to extend validation beyond a single benchmarked dataset, we developed ECG-Note using publicly available datasets and large language models, establishing an additional external validation dataset for LEF detection that captures diverse patient populations and clinical contexts.
@@ 4.3 Moving beyond end-to-end black-box modeling
Most AI-ECG approaches for left ventricular dysfunction have relied on end-to-end deep learning architectures (Attia et al., 2019; Attia and others, 2019; Carter et al., 2026; Poterucha et al., 2025). While such models achieve strong predictive performance, their limited interpretability poses challenges for clinical integration, particularly in applications that extend beyond conventional ECG criteria grounded in established ECG principles. In the absence of transparent mechanistic reasoning, black-box predictions may be difficult to reconcile with established diagnostic frameworks, potentially limiting clinician trust and adoption. Efforts to enhance transparency have included tabular models constructed from 555 discrete ECG measurements (Hughes et al., 2024). Although more interpretable, these approaches depend on proprietary commercial measurement algorithms that may vary across ECG platforms (Strodthoff et al., 2023), thereby constraining generalizability and standardized external evaluation. In contrast, we introduce a structured representation paradigm that integrates the predictive capacity of deep learning with the interpretability of tabular modeling. Each predictor corresponds to clinically meaningful features and it does not rely on vendor-specific measurement pipelines. Within the publicly benchmarked setting, this approach demonstrates improved performance over the end-to-end Columbia mini model, supporting the feasibility of clinically interpretable yet high-performing AI-ECG systems.
@@ A.2 Construction of the external test set
We constructed the external test set ECG-Note as Figure A.1. Adult patients, who had at least one standard 10 s 12-lead ECG and at least one clinical note within one year following the ECG from MIMIC-IV (Johnson et al., 2024), MIMIC-IV-ECG (Gow et al., 2023), and MIMIC-IV-Note (Johnson et al., 2023b), were initially considered, leading to a starting pool of 1,113,547 ECG- note pairs from 103,505 patients with 521,654 ECGs and 247,313 notes. Exclusion criteria included: 1) ECG-Note pairs where the clinical note did not contain keywords related to ejection fraction(EF) and Echo/TTE; 2) ECG-Note pairs where the ECG contained NaN data. After exclusions, 382,308 ECG-Note paired data of 37,234 patients, comprising 246,543 ECGs and 62,023 notes, were included for further analysis. For the note set, we employed a Large Language Model (LLM), specifically Qwen2.5-72B-Instruct(Qwen et al., 2025), to extract precise ejection fraction(EF) values derived solely from the concurrent Echo/TTE (Fig.A.1 b). Based on the values obtained by the LLM, we determined whether the ejection fraction was $\leq$ 45%, suggesting moderately reduced systolic function. These labels were cross-referenced with MIMIC-ECG-LVEF (Li et al., 2025) to identify consistent ECG-EF class pairs, resulting in the collection of 55,318 available ECGs. To ensure the accuracy of subsequent evaluations, only the ECG-Note pair with the shortest time interval for each patient was retained among the valid samples, resulting in a final testing set of 16,017 ECGs.
@@ B.2 Subgroup analysis for multi-predictor method in the external test set
To evaluate model robustness and fairness across patient characteristics and clinical settings, we conducted a subgroup analysis comparing our proposed Transformer-PT-XGBoost with Columbia mini model. Subgroups were defined by age, sex, race and ethnicity, and clinical context. Data linkage between the MIMIC-IV and MIMIC-IV-ECG datasets was performed to extract these subgroup attributes. Patient age was recalculated specifically at the time of ECG acquisition by leveraging the time-invariant interval method to restore temporal alignment (Johnson et al., 2023a). For race and ethnicity, we aggregated granular categories into broader clusters (Table B.3). Regarding clinical context, we determined the admission type by mapping the ECG timestamp to the corresponding hospital admission and discharge window. All performance metrics are reported using the AUROC and the AUPRC (Table B.2).

### 05 Apr 2026 v2

@@ 1 Introduction
Heart failure (HF) is a major global health burden, affecting over 64 million individuals worldwide, with prevalence continuing to rise (Savarese et al., 2022). In many patients, left ventricular systolic dysfunction (LVSD) is initially asymptomatic, yet it frequently progresses to symptomatic HF and is associated with poor prognosis if left undetected. Low left ventricular ejection fraction (LEF) is a critical indicator that reflects the progression of LVSD toward clinically manifest HF (Heidenreich et al., 2022). Early detection of LVSD is critical, as timely initiation of evidence-based therapies can mitigate symptoms, slow disease progression, and improve survival. Although cardiovascular imaging, particularly echocardiography (ECHO), remains the clinical gold standard for assessing ventricular function, it is time- and resource-intensive, requires specialized operators and equipment, and is not readily scalable for population-level screening (Ciampi and Villari, 2007). These limitations underscore the need for cost-effective, widely accessible approaches capable of reliably identifying patients with LEF (Attia et al., 2019a; Yao et al., 2021; Tran et al., 2025; Lakshmanan and Mbanze, 2023).
@@
Electrocardiography (ECG), which records cardiac electrical activity, is a promising tool for large-scale screening (Hughes et al., 2024). Since its introduction by Willem Einthoven, which earned him the Nobel Prize in 1924, ECG science has steadily advanced: acquisition has become easier, signal quality has improved, and pathognomonic features for a wide range of rhythm and conduction disorders have been defined (Khera, 2024). More recently, artificial intelligence (AI) has transformed ECG analysis, enabling models to achieve cardiologist-level performance in traditional interpretation tasks (Hannun et al., 2019; Ribeiro et al., 2020; Jiang et al., 2024; Li et al., 2025) and to detect subtle patterns associated with conditions such as aortic stenosis (Kwon et al., 2020), tricuspid regurgitation (Diao et al., 2025), and LEF (Attia et al., 2019a), which are often imperceptible to humans. Despite these advances, clinical adoption varies: AI-ECG models are widely accepted for well-established conditions such as atrial fibrillation and sinus bradycardia, which can be visually confirmed according to guidelines (Joglar et al., 2024; Kusumoto et al., 2019), but remain limited for emerging tasks such as LEF detection due to concerns about interpretability and the difficulty of linking predictions to recognizable ECG features (Van De Leur et al., 2022; Hughes et al., 2024). To enhance interpretability and facilitate broader clinical deployment, a recent study proposed a random forest model based on 555 discrete ECG measurements based on commercial algorithms, which demonstrated superior performance compared with N-terminal prohormone brain natriuretic peptide (NT-proBNP) for LEF detection (Hughes et al., 2024). Although this approach showed promising results, its performance is slightly below deep learning. Moreover, as highlighted in the original study, many features cannot be obtained when applied across different ECG machines, and measurement values differ markedly between devices (Strodthoff et al., 2023), which presents a significant barrier to broader clinical adoption.
@@
In this study, we aim to evaluate the efficacy of interpretable methods for detecting LEF using standard 12-lead ECGs. We propose ECG-based Predictor-Driven LEF (ECGPD-LEF), a framework that integrates traditional AI-ECG interpretation models with either a single-predictor approach or a multi-predictor approach to enhance recognition of subtle ECG features and improve LEF detection. The single-predictor approach performs zero-shot-like inference, requiring no additional training on the target dataset. The multi-predictor approach was trained on the publicly available ECG-ECHO paired dataset EchoNext, a comprehensive dataset encompassing diverse populations across races and clinical contexts. We also constructed an external validation dataset, MIMIC-LEF, based on MIMIC-IV (Johnson et al., 2024), MIMIC-IV-ECG (Gow et al., 2023), and MIMIC-IV-Note(Johnson et al., 2023a), which similarly covers a heterogeneous population. The model was evaluated on multi-center datasets, including internal validation (the EchoNext test set) and external validation (MIMIC-LEF). The model performance was assessed across multiple dimensions, including the presence or absence of other structural heart disease (SHD) and valvular heart disease (VHD), as well as across different care settings and racial or ethnic subgroups. Global and local model explanations were performed to identify important predictors contributing to LEF detection, and to characterize these subtle ECG changes associated with LEF.
@@ 2.1 Data sources
For the detection of LEF, we used the ECG-ECHO paired EchoNext dataset collected at Columbia University Irving Medical Center (Elias and Finer, 2025). External validation was performed using an ECG-Note paired dataset constructed from MIMIC-IV (Johnson et al., 2024), MIMIC-IV-ECG (Gow et al., 2023), and MIMIC-IV-Note (Johnson et al., 2023a), comprising records from Beth Israel Deaconess Medical Center.
@@ 2.3 ECG-ECHO and ECG-Note datasets
The ECG-Note dataset is constructed based on MIMIC-IV (Johnson et al., 2024), MIMIC-IV-ECG (Gow et al., 2023), and MIMIC-IV-Note (Johnson et al., 2023a), as illustrated in Figure A.1 of the Appendix. We initially identified adult patients who possessed at least one standard 10-second 12-lead ECG with sampling frequency of 500Hz and a corresponding clinical note recorded within one year following the ECG. Exclusion criteria were applied to remove: 1) pairs where clinical notes lacked keywords related to ejection fraction(EF) and ECHO/TTE; 2) ECGs containing missing data (NaN). Following these exclusions, the samples were cross-referenced with MIMIC-ECG-LVEF (Li et al., 2025) to identify consistent ECG-EF class pairs. For each patient, only the ECG-Note pair with the shortest time interval was retained, resulting in a final testing set of 16,017 ECGs.
@@ 2.7 Interpretability
For the multi-predictor approach, we employed SHAP (SHapley Additive exPlanations) values (Shapley et al., 1953; Lundberg and Lee, 2017) to quantify feature contributions at global and local levels. SHAP is a game-theoretic, additive feature attribution method that provides consistent and locally accurate explanations (Lundberg and Lee, 2017). For computational efficiency in tree-based models, we used the Tree SHAP algorithm (Lundberg et al., 2018). Global explanation plots included cumulative contributions, beeswarm summaries, and SHAP versus predictor value plots to identify key predictors and characterize their behavior in the model. Local plots highlighted the ten predictors SHAP values and displayed predicted probabilities relative to both the ECG diagnosis thresholds and the LEF decision thresholds, facilitating interpretation of individual model predictions.
@@ 3.1 Population characteristics
  • The training, validation, and internal test splits correspond to the official splits of the ECG-ECHO dataset, EchoNext (Elias and Finer, 2025). The external test cohort, ECG-Note, was derived from the MIMIC-IV database and its associated ECG and clinical note modules (Johnson et al., 2024; Gow et al., 2023; Johnson et al., 2023a). Values are shown as counts and percentages.
@@ 4.2 The need for publicly benchmarked AI-ECG evaluation
Artificial intelligence applied to ECG has demonstrated strong performance in emerging tasks, including detection of aortic stenosis, tricuspid regurgitation and left ventricular dysfunction. However, most prior models have been developed and evaluated on institution-specific private datasets with heterogeneous population characteristics and outcome definitions, limiting reproducibility and preventing rigorous cross-model comparison. Early work by Attia et al. (2019a) trained a convolutional neural network on 44,959 patients to identify ventricular dysfunction defined as ejection fraction $\leq$ 35%, with prospective validation performed at the same institution (Attia et al., 2019b). Subsequent refinements extended detection to ejection fraction $\leq$ 40% and reported multisite validation using digital ECG input alone (Carter et al., 2026). Although these studies achieved strong discrimination, the absence of publicly accessible benchmarking datasets constrains transparent evaluation. Recently, Poterucha et al. (2025) addressed this gap by releasing a de-identified ECG dataset comprising 36,286 unique patients with predefined training, validation, and test splits, and by benchmarking a baseline model (the Columbia mini model) that achieved discrimination comparable to models trained on substantially larger proprietary cohorts. By providing standardized data partitions, model weights, and open-source code, this work established a transparent and reproducible evaluation framework for AI-ECG research, upon which our study enables rigorous and fair comparison. Furthermore, to extend validation beyond a single benchmarked dataset, we developed ECG-Note using publicly available datasets and large language models, establishing an additional external validation dataset for LEF detection that captures diverse patient populations and clinical contexts.
@@ 4.3 Moving beyond end-to-end black-box modeling
Most AI-ECG approaches for left ventricular dysfunction have relied on end-to-end deep learning architectures (Attia et al., 2019a, b; Carter et al., 2026; Poterucha et al., 2025). While such models achieve strong predictive performance, their limited interpretability poses challenges for clinical integration, particularly in applications that extend beyond conventional ECG criteria grounded in established ECG principles. In the absence of transparent mechanistic reasoning, black-box predictions may be difficult to reconcile with established diagnostic frameworks, potentially limiting clinician trust and adoption. Efforts to enhance transparency have included tabular models constructed from 555 discrete ECG measurements (Hughes et al., 2024). Although more interpretable, these approaches depend on proprietary commercial measurement algorithms that may vary across ECG platforms (Strodthoff et al., 2023), thereby constraining generalizability and standardized external evaluation. In contrast, we introduce a structured representation paradigm that integrates the predictive capacity of deep learning with the interpretability of tabular modeling. Each predictor corresponds to clinically meaningful features and it does not rely on vendor-specific measurement pipelines. Within the publicly benchmarked setting, this approach demonstrates improved performance over the end-to-end Columbia mini model, supporting the feasibility of clinically interpretable yet high-performing AI-ECG systems.
@@ A.2 Construction of the external test set
We constructed the external test set ECG-Note as Figure A.1. Adult patients, who had at least one standard 10 s 12-lead ECG and at least one clinical note within one year following the ECG from MIMIC-IV (Johnson et al., 2024), MIMIC-IV-ECG (Gow et al., 2023), and MIMIC-IV-Note (Johnson et al., 2023a), were initially considered, leading to a starting pool of 1,113,547 ECG- note pairs from 103,505 patients with 521,654 ECGs and 247,313 notes. Exclusion criteria included: 1) ECG-Note pairs where the clinical note did not contain keywords related to ejection fraction(EF) and Echo/TTE; 2) ECG-Note pairs where the ECG contained NaN data. After exclusions, 382,308 ECG-Note paired data of 37,234 patients, comprising 246,543 ECGs and 62,023 notes, were included for further analysis. For the note set, we employed a Large Language Model (LLM), specifically Qwen2.5-72B-Instruct(Qwen et al., 2025), to extract precise ejection fraction(EF) values derived solely from the concurrent Echo/TTE (Fig.A.1 b). Based on the values obtained by the LLM, we determined whether the ejection fraction was $\leq$ 45%, suggesting moderately reduced systolic function. These labels were cross-referenced with MIMIC-ECG-LVEF (Li et al., 2025) to identify consistent ECG-EF class pairs, resulting in the collection of 55,318 available ECGs. To ensure the accuracy of subsequent evaluations, only the ECG-Note pair with the shortest time interval for each patient was retained among the valid samples, resulting in a final testing set of 16,017 ECGs.
@@ B.2 Subgroup analysis for multi-predictor method in the external test set
To evaluate model robustness and fairness across patient characteristics and clinical settings, we conducted a subgroup analysis comparing our proposed Transformer-PT-XGBoost with Columbia mini model. Subgroups were defined by age, sex, race and ethnicity, and clinical context. Data linkage between the MIMIC-IV and MIMIC-IV-ECG datasets was performed to extract these subgroup attributes. Patient age was recalculated specifically at the time of ECG acquisition by leveraging the time-invariant interval method to restore temporal alignment (Johnson et al., 2023b). For race and ethnicity, we aggregated granular categories into broader clusters (Table B.3). Regarding clinical context, we determined the admission type by mapping the ECG timestamp to the corresponding hospital admission and discharge window. All performance metrics are reported using the AUROC and the AUPRC (Table B.2).